# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ 3.7 Bound on $T_{6}$
$|T_{6,HL,L,HL}^{x,1}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
For $T_{6,HL,L,HL}^{x,2}$, still using Cauchy-Schwarz inequality in $y$ and Young’s convolution inequality,
$|T_{6,HL,L,HL}^{x,2}|$ $\lesssim\iint 1_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}|k|^{\frac{1}{2}}\|\psi_{k}\|_{\infty}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}|k^{\prime}|^{-\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\Delta_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k.$
By Gagliardo-Nirenberg-Sobolev inequality, the first term obeys
$\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{2}\|\partial_{y}\psi_{k}\|_{2}\mathrm{d}k\lesssim\nu^{-\frac{2}{3}}\mathcal{E}.$
For the third term, it is the same as the third term of $|T_{6,HL,L,HL}^{x,1}|$, so
$\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$
Combining the three parts, we have
$|T_{6,HL,L,HL}^{x,2}|\lesssim(\nu^{-\frac{2}{3}}\mathcal{E})^{\frac{1}{2}}(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}=\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
The estimate of $T_{6,HL,L,HL}^{y}$ is similar to $T_{6,HL,L,HL}^{x,1}$, and then we have $|T_{6,HL,L,HL}^{y}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}$.
3.8 Bound on $T_{7}$
Bounding the term is the simplest. We just use Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4,
$|T_{7}|\lesssim\sup_{k}\int\|y\omega_{k}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\lesssim\sup_{k}\|y\omega_{k}\|_{2}\sup_{k}\|y\nabla_{k}\omega_{k}\|_{2}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@ 3.9 Bound on $T_{8}$
By integration by parts,
$T_{8}=-2\sup_{k}\mathrm{Re}\langle\psi_{k},NL_{k}\rangle.$
We separate it into two parts,
@@
$T_{8}=2\sup_{k}\mathrm{Re}\left\langle\psi_{k},\int i(k-k^{\prime})\psi_{k-k^{\prime}}\partial_{y}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle-2\sup_{k}\mathrm{Re}\left\langle\psi_{k},\int\partial_{y}\psi_{k-k^{\prime}}ik^{\prime}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\eqqcolon T_{8}^{x}+T_{8}^{y}.$
For $T_{8}^{x}$, we apply integration by parts and separate it into two parts again,
@@
$T_{8}^{x}=-2\sup_{k}\int\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k^{\prime}-2\sup_{k}\int\mathrm{Re}\langle\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k^{\prime}\eqqcolon T_{8}^{x,1}+T_{8}^{x,2}.$
By Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4, $T_{8}^{x,1}$ obeys
$|T_{8}^{x,1}|\lesssim\sup_{k}\int\|\partial_{y}\psi_{k}\|_{2}\|(k-k^{\prime})\psi_{k-k^{\prime}}\|_{\infty}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\leq\sup_{k}\|\nabla_{k}\psi_{k}\|_{2}\sup_{k}\|\omega_{k}\|_{2}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
Turning to $T_{8}^{x,2}$, we similarly have
@@
$|T_{8}^{x,2}|\lesssim\sup_{k}\int\|\psi_{k}\|_{\infty}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\lesssim\left(\sup_{k}\|\omega_{k}\|_{2}\right)^{2}\int\|\psi_{k}\|_{\infty}\mathrm{d}k,$
where we have used $\|k\partial_{y}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2}$. For the last term, using Gagliardo-Nirenberg-Sobolev inequality,
$\int\|\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\int|k|^{-\frac{1}{2}}|k|^{\frac{1}{2}}\|\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\lesssim\int|k|^{-\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k.$
To estimate the term, we split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|\leq\nu^{-\frac{1}{3}}$, then the $|k|\geq\nu^{-\frac{1}{3}}$ term obeys
$\int_{|k|\geq\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}(\nu^{-\frac{1}{2}}|k|^{\frac{1}{2}})^{-1}|k|^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\nu^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}.$
Then we integrate over $|k|\leq\nu^{-\frac{1}{3}}$,
$\int_{|k|\leq\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\lesssim\int_{|k|\leq\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\mathrm{d}k\sup_{k}\|\nabla_{k}\psi_{k}\|_{2}\lesssim\nu^{-\frac{1}{6}}\mathcal{E}^{\frac{1}{2}}.$
Combining the two parts, we have
$|T_{8}^{x,2}|\lesssim\nu^{-1}\mathcal{D}\nu^{-\frac{1}{6}}\mathcal{E}^{\frac{1}{2}}=\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
Turning to $T_{8}^{y}$, we use the decomposition $1=1_{|k-k^{\prime}|\leq\frac{|k|}{2}}+1_{|k-k^{\prime}|>\frac{|k|}{2}}$, and hence we write $T_{8}^{y}\eqqcolon T_{8,LH}^{y}+T_{8,HL}^{y}$. For $T_{8,LH}^{y}$, using $|k^{\prime}|\approx|k|$, the estimate is similar to $T_{8}^{x,1}$,
$|T_{8,LH}^{y}|\lesssim\sup_{k}\int|k|\|\psi_{k}\|_{2}\|\partial_{y}\psi_{k-k^{\prime}}\|_{\infty}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\lesssim\sup_{k}\|\nabla_{k}\psi_{k}\|_{2}\sup_{k}\|\omega_{k}\|_{2}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
While for $T_{8,HL}^{y}$, using $|k^{\prime}|\lesssim|k-k^{\prime}|$, the estimate is similar to $T_{8}^{x,2}$,
@@
$|T_{8,HL}^{y}|\lesssim\sup_{k}\int\|\psi_{k}\|_{\infty}|k-k^{\prime}|\|\partial_{y}\psi_{k-k^{\prime}}\|_{2}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\lesssim\left(\sup_{k}\|\omega_{k}\|_{2}\right)^{2}\int\|\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
3.10 Bound on $\mathcal{NL}$
@@
$|\mathcal{NL}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}\leq-4c\mathcal{D}+C\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
Thus if $\mathcal{E}(0)\leq c^{2}C^{-\frac{1}{2}}\nu^{\frac{7}{12}}$, then
$\sup_{t\in[0,+\infty)}\mathcal{E}(t)+2c\int_{0}^{+\infty}\mathcal{D}(s)\mathrm{d}s\leq\mathcal{E}(0).$
And then Theorem 1.1 follows directly from Theorem 3.6.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China grants 11571333, 11471301. The author would like to thank the advisor for his guidance Lifeng Zhao.

### v4

@@ 3.6 Bound on $T_{7}$
$|T_{7}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@ 3.7 Bound on $T_{8}$
We handle $T_{8}$ using integration by parts,
$T_{8}$ $=4\sup_{k\in\mathbb{R}}\int_{0}^{t}-\mathrm{Re}\langle\psi_{k},NL_{k}\rangle\mathrm{d}s$
@@
$=4\sup_{k\in\mathbb{R}}\int_{0}^{t}\mathrm{Re}\left\langle\psi_{k},\int_{\mathbb{R}}i(k-k^{\prime})\psi_{k-k^{\prime}}\partial_{y}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle-\mathrm{Re}\left\langle\psi_{k},\int_{\mathbb{R}}\partial_{y}\psi_{k-k^{\prime}}ik^{\prime}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\mathrm{d}s$
@@
$=4\sup_{k\in\mathbb{R}}\int_{0}^{t}\int_{\mathbb{R}}-\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}}+\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle-\mathrm{Re}\left\langle\psi_{k},\partial_{y}\psi_{k-k^{\prime}}ik^{\prime}\omega_{k^{\prime}}\right\rangle\mathrm{d}k^{\prime}\mathrm{d}s$
@@
$=4\sup_{k\in\mathbb{R}}\int_{0}^{t}\int_{\mathbb{R}}-\mathrm{Re}\langle\partial_{y}\psi_{k},i(k-k^{\prime})\psi_{k-k^{\prime}}\omega_{k^{\prime}}\rangle-\mathrm{Re}\left\langle k\psi_{k},i\partial_{y}\psi_{k-k^{\prime}}\omega_{k^{\prime}}\right\rangle\mathrm{d}k^{\prime}\mathrm{d}s.$
By Cauchy-Schwarz inequality in $y$, Hölder’s inequality in $s$, Young’s convolution inequality and Lemma 3.5, we have
$|T_{8}|$ $\lesssim\sup_{k\in\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{t}\|\nabla_{k}\psi_{k}\|_{2}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}s\mathrm{d}k^{\prime}$
@@
$\lesssim\sup_{k\in\mathbb{R}}\int_{\mathbb{R}}\sup_{s\in[0,t]}\|\nabla_{k}\psi_{k}\|_{2}\left(\int_{0}^{t}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\|\omega_{k^{\prime}}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k^{\prime}$
$\lesssim\sup_{k\in\mathbb{R}}\sup_{s\in[0,t]}\|\nabla_{k}\psi_{k}\|_{2}\cdot\int_{\mathbb{R}}\left(\int_{0}^{t}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\|\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
3.8 Bound on $\mathbb{NL}$
@@
$|\mathbb{NL}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$\mathcal{E}(t)\leq 2\mathcal{E}(0)-4c\mathcal{D}(t)+C\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$ (3.5)
Thus if $\mathcal{E}(0)\leq c^{2}C^{-2}\nu^{\frac{7}{3}}$, then
$\sup_{t\in[0,+\infty)}\mathcal{E}(t)\leq 2\mathcal{E}(0).$ (3.6)
Proof.
(3.5) follows directly from (3.2) and the estimate of $\mathbb{NL}$. Now we suppose that $\mathcal{E}(0)\leq c^{2}C^{-2}\nu^{\frac{7}{3}}$ and $\mathcal{E}(0)\neq 0$, then we demonstrate (3.6) by contradiction. Assume that (3.6) does not hold, that is, there exists $t_{0}\in[0,+\infty)$ such that $\mathcal{E}(t_{0})>2\mathcal{E}(0)$. Let
$t^{*}=\inf\{t\geq 0|\mathcal{E}(t)=\min\{4\mathcal{E}(0),\mathcal{E}(t_{0})\}\}\leq t_{0}<\infty,$
then $\mathcal{E}(t)\leq\min\{4\mathcal{E}(0),\mathcal{E}(t_{0})\}\leq 4\mathcal{E}(0)$ for any $t\in[0,t^{*}]$ and $\mathcal{E}(t^{*})=\min\{4\mathcal{E}(0),\mathcal{E}(t_{0})\}>2\mathcal{E}(0)$. However, according to (3.5),
$\mathcal{E}(t^{*})\leq 2\mathcal{E}(0)+\left(-4c+C\nu^{-\frac{7}{6}}(4\mathcal{E}(0))^{\frac{1}{2}}\right)\mathcal{D}(t)\leq 2\mathcal{E}(0),$
which leads to a contradiction. ∎
Theorem 1.1 follows directly from Theorem 3.7.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China 12271497. The author would like to thank professor Lifeng Zhao for his guidance.